# Multi-scale Cascaded Large-Model for Whole-body ROI Segmentation

### v1

@@ IV-D6 Experiments on KiTS (2023)
This dataset builds upon the segmentation labels of kidneys and kidney tumors from KiTS (2019) by adding segmentation for cysts. As shown in Table VI, traditional networks do not perform well in the segmentation of tumors and cysts. For example, the best-performing TransUNet achieves an average DSC of only 42.31% for tumors and 42.68% for cysts. In contrast, our designed model improves the DSC for cyst segmentation to 55.31%, and MCFNet achieves the best performance in segmentation for kidneys, kidney tumors, and cysts.
@@
As shown in the sixth row of Fig. 7, the addition of cyst segmentation labels increases the difficulty of segmenting small target areas. Models like UNet, TransUNet, and Swin-Unet demonstrate shortcomings in segmenting cysts effectively. Additionally, while UCTransNet accurately identifies cyst regions, it struggles to maintain accurate recognition of tumor regions simultaneously.
IV-D7 Experiments on LUNG
In this dataset, we select four evaluation metrics to validate the model’s segmentation performance, as shown in Table VII. It is observed that the advanced model FSCA-Net achieves the best results in terms of DSC and Recall, while Cascaded MERIT achieves the best Precision among existing models, reaching 82.23%. Our designed model achieves the best performance across DSC, HD95, Recall, and Precision metrics. Specifically, MCFNet improves DSC by a significant 2.39% and Recall by 2.99% compared to FSCA-Net.
As shown in the seventh row of Fig. 7, for single-label segmentation targets, most models correctly identify tumor regions except for Swin-Unet. UNet identifies tumor regions but also mistakenly includes background areas as tumors. Swin-Unet not only fails to accurately identify tumor regions but also misidentifies vascular regions in the background as tumors.
@@ IV-D8 Experiments on CPCGEA
As shown in Table VIII, we compare seven comparison models for the segmentation performance of prostate cancer. Since we are dealing with single-label image segmentation, we use four evaluation metrics to comprehensively assess model performance. Our designed model improves upon the current best model by 1.78% in DSC, reduces HD95 by 0.582mm, increases Recall by 1.01%, and improves Precision by 0.29%.
As shown in the eighth row of Fig. 7, we continue the work initiated by our previously proposed model, FSCA-Net, and conducted further exploration on CPCGEA. We add experiments with TransCASCADE and Cascaded MERIT, revealing that they exhibit unclear boundary segmentation for prostate cancer regions and struggle to delineate the edges of tumor areas effectively.
TABLE XI: ABLATION EXPERIMENTS ON FIVE DATASETS.
@@
Figure 9: The experiments shown in the figure are conducted on the CPCGEA dataset. (a) Visualization of model performance and parameter size. (b) Visualization of model performance and FLOPs.
IV-D9 Experiments on KNEE
As shown in Table IX, all models achieve an average DSC for femur segmentation exceeding 97.5%, with our designed model achieving the highest at 99.14%. Additionally, HD95, Recall, and Precision also achieve optimal performance, although the improvement in segmentation performance compared to advanced models is relatively limited.
As depicted in the ninth row of Fig. 7, we can see that all models accurately segment the femur region. However, Swin-Unet and UCTransNet exhibit minor segmentation errors in very small areas, although the overall differences are minimal.
IV-D10 Experiments on AutoPET
For whole-body tumor segmentation, our designed model and seven comparison models are evaluated using four metrics, as shown in the Table X. Among the current advanced methods, TransUNet achieves the best DSC of 61.95%, the best HD95 of 22.329mm, and the best Precision of 72.98%. Additionally, FSCA-Net achieves a Recall of 62.07%. However, our designed model MCFNet improves by 1.92% in DSC, reduces HD95 by 0.078mm, increases Recall by 3.28%, and improves Precision by 0.93%.
As shown in Fig. 8, four segmentation cases from the AutoPET dataset are presented. The erroneous segmentation areas are all marked with red boxes. In the whole-body tumor segmentation images, there are many background areas that appear white and have smooth edges but are not tumors, leading the segmentation models to mistakenly identify tumor regions. From a comprehensive observation of the four cases, it can be seen that most models exhibit confusion between the background and tumor regions. Specifically, UNet and TransUNet have issues with over-segmentation and under-segmentation. Swin-Unet, UCTransNet, and FSCA-Net all display errors in segmentation, while TransCASCADE and Cascaded MERIT show cases of under-segmentation. Although the MCFNet model has performed relatively well in tumor segmentation, there is still room for improvement in segmenting the detailed boundaries of tumor regions.
@@ IV-F Discussion
MCFNet integrates low parameter count, low computational cost, and high segmentation performance. Our designed model effectively balances performance with parameter and computational cost, demonstrating significant advantages over existing state-of-the-art methods. To illustrate this, we conduct experiments on the CPCGEA dataset and compare our model’s results with several benchmark models. These experimental results are visualized in figures, clearly showing the performance of the MCFNet model. As shown in Fig. 9, our model not only exhibits excellent segmentation performance but also achieves or surpasses the performance of existing methods while maintaining a smaller parameter size and lower computational cost. This indicates that our model offers higher efficiency and better scalability in practical applications. The optimization of performance, computational cost, and parameter size makes our model suitable for resource-constrained environments and lays a solid foundation for its future promotion and application in various scenarios.
@@ V Conclusion
We design a novel multi-scale, multi-resolution cascade fusion network called MCFNet. By capturing rich image features from multi-scale and multi-resolution images, our designed model is better suited to various application scenarios. We propose an adaptive loss aggregation strategy called Adaptive-MFA to optimize model training and performance. We conduct experiments on multiple datasets covering whole-body risk organs and tumors to fully validate the robustness and generalization capability of the model. Experimental results show that our MCFNet exhibits excellent segmentation performance and good generalization ability across multiple datasets.
In practical clinical applications, the efficient segmentation capability of MCFNet can significantly enhance the accuracy and efficiency of medical image analysis, reduce the workload of doctors, and improve the precision of diagnosis and treatment. By accurately segmenting risk organs and tumor regions, MCFNet can assist doctors in better formulating treatment plans, thereby improving patient outcomes. Additionally, the good generalization ability of MCFNet means it can adapt to image data generated by different hospitals and equipment, making it a promising tool for widespread use in clinical practice. This can drive advancements and progress in the field of medical image analysis.

### v2

@@ IV-D6 Experiments on KiTS (2023)
As shown in Table VI, traditional networks perform poorly in tumor and cyst segmentation, with the best TransUNet achieving a DSC of only 42.31% for tumors and 42.68% for cysts. In contrast, our model improves the DSC for cyst segmentation to 55.31%, and MCFNet excels in segmenting kidneys, tumors, and cysts.
As shown in the second row of Fig. 9, cyst segmentation is more challenging due to small target areas. Models like TransUNet, and Swin-Unet struggle with cyst segmentation. While UCTransNet correctly identifies cyst regions, it has difficulty maintaining accurate tumor segmentation simultaneously.
@@
IV-D7 Experiments on LUNG
We evaluate segmentation performance using four metrics, as summarized in Table VII. Among the baselines, FSCA-Net achieves the best DSC and Recall, while Cascaded MERIT excels in Precision at 82.23%. Our model outperforms all models in DSC, HD95, Recall, and Precision. Specifically, MCFNet improves DSC by 2.39% and Recall by 2.99% compared to FSCA-Net.
As shown in the third row of Fig. 9, most models correctly identify tumor regions, except Swin-Unet. UNet includes background areas as tumors, while Swin-Unet misidentifies both tumor and vascular regions as tumors.
Figure 10: Visual examples of segmentation on the AutoPET dataset. The tumor lesion region is marked in red for clarity.
Figure 11: The visual results of our method on the AutoPET dataset. The tumor lesion region is marked in red for clarity.
@@ IV-D8 Experiments on CPCGEA
As shown in Table VIII, we evaluate seven models for prostate cancer segmentation using four metrics. The proposed method attains the best overall performance, exceeding the previous state of the art by 1.78% in DSC, reducing HD95 by 0.582 mm, and yielding gains of 1.01% in Recall and 0.29% in Precision.
As shown in the fourth row of Fig. 9, the comparative models fail to clearly delineate the boundaries of the tumor region. TransCASCADE and Cascaded MERIT are prone to mistakenly identifying background areas as lesions.
IV-D9 Experiments on KNEE
As shown in Table IX, all models achieve an average DSC for femur segmentation exceeding 97.5%, with MCFNet achieving the highest at 99.14%. Furthermore, HD95, Recall, and Precision also achieve optimal performance, although the improvement in segmentation performance compared to advanced models is relatively limited.
As depicted in the fifth row of Fig. 9, we can see that all models accurately segment the femur region. However, Swin-Unet and UCTransNet exhibit minor segmentation errors in very small areas, although the overall differences are minimal.
IV-D10 Experiments on AutoPET
For whole-body tumor segmentation, our designed model and seven comparison models are evaluated using four metrics, as shown in the Table X. TransUNet achieves the best DSC of 61.95%, the best HD95 of 22.329mm, and the best Precision of 72.98%. Additionally, FSCA-Net achieves a Recall of 62.07%. However, our designed model MCFNet improves by 1.92% in DSC, reduces HD95 by 0.078mm, increases Recall by 3.28%, and improves Precision by 0.93%.
Fig. 10 and Fig. 11 present four segmentation cases from the AutoPET dataset, with erroneous areas marked in red boxes. In the whole-body tumor segmentation images, there are many background areas that appear white and have smooth edges but are not tumors, leading the segmentation models to mistakenly identify tumor regions. Most models confuse background with tumor regions, with UNet and TransUNet showing over- and under-segmentation. Swin-Unet, UCTransNet, and FSCA-Net also exhibit errors, while TransCASCADE and Cascaded MERIT show under-segmentation. MCFNet performs well but can still improve in segmenting detailed tumor boundaries.
TABLE XI: ABLATION EXPERIMENTS ON SEGTHOR, CHAOS, LITS, KITS2019, AND CPCGEA DATASETS.
@@ IV-D9 Experiments on KNEE
TABLE XII: COMPLEXITY COMPARISON BETWEEN MLP AND LINEAR LAYERS IN THE LAT MODULE.
[…]
(G) (M)
LAT w/ MLP 61.2 57.7
LAT w/ Linear 59.3 47.0
@@ IV-F Discussion
IV-F1 Discussion on the Complexity Optimization of the LAT Module
As shown in Table XII, replacing the MLP layer with a linear layer in the LAT module effectively reduces model complexity. Specifically, the number of parameters decreases from 57.7M to 47.0M, and the computational cost (FLOPs) is reduced from 61.2G to 59.3G. The structural simplification is achieved by removing redundant non-linear transformations present in the MLP. Consequently, this optimization yields a more efficient and practical LAT module, making it particularly suitable for resource-constrained application scenarios.
TABLE XIII: ACCURACY (DICE SCORE) VS. MODEL COMPLEXITY (PARAMETERS AND COMPUTATIONAL COMPLEXITY) COMPARISON ON THE CPCGEA DATASET.
[…]
(%, mean) (G) (M)
U-Net 64.1 50.3 14.8
TransUNet 63.2 56.7 105.0
SwinUnet 60.9 67.3 82.3
UCTransNet 60.3 63.2 65.6
TransCASCADE 65.4 22.1 123.5
Cascaded MERIT 65.5 33.3 147.9
FSCA-Net 65.2 32.7 44.2
Ours 67.3 59.3 47.0
IV-F2 Discussion on the Model Complexity
MCFNet integrates low parameter count, low computational cost, and high segmentation performance. Our designed model effectively balances performance with parameter and computational cost, demonstrating significant advantages over existing state-of-the-art methods. To substantiate the approach, we evaluate on the CPCGEA dataset, comparing our model with several strong baselines. Figure 2 and Table XIII clearly indicate that MCFNet achieves excellent segmentation performance alongside lower parameter count and computation than competing approaches. These results indicate that our model delivers higher computational efficiency and superior scalability in real-world deployments. The optimization of performance, computational cost, and parameter size makes our model suitable for resource-constrained environments and lays a solid foundation for its future application in various scenarios.
@@ V Conclusion
We design a novel multi-scale cascaded fusion network called MCFNet. This network cascades two backbones, FCB and SEB, to capture rich features from multi-scale and multi-resolution images. Among them, we introduce an innovative Linear Attention Transformer module that ensures efficient feature extraction while reducing computational and storage requirements. Finally, we propose an adaptive loss aggregation strategy called Adaptive-MFA to optimize model training and performance. Extensive experiments on multiple benchmarks demonstrate that MCFNet achieves high segmentation performance with strong generalization.
MCFNet delivers efficient, high-accuracy segmentation, improving medical image analysis and enhancing clinical precision. By accurately segmenting OARs and tumor regions, it assists doctors in formulating better treatment plans, thereby improving patient outcomes. Its strong generalization ability allows MCFNet to adapt to data generated by different hospitals and equipment, making it a promising tool for widespread clinical application and advancing medical image analysis.